OpenStudy (anonymous):

Simplify -14x^3/ x^3 - 5x^4

OpenStudy (anonymous):

this is the same as the last question you asked. the steps are the same

OpenStudy (anonymous):

would x not equai 1/5 and 0 or just 1/5 ?

OpenStudy (anonymous):

you're trying to simplify yes? not solve for x

OpenStudy (anonymous):

but i also need to find what x doesnt equal

OpenStudy (anonymous):

what happens if x = 0?

OpenStudy (anonymous):

then theres no solution

OpenStudy (anonymous):

why is there no solution?

OpenStudy (anonymous):

looks like there is a solution to me

OpenStudy (anonymous):

then is it 0?

OpenStudy (anonymous):

how does it show that it cant be 1/5?

OpenStudy (anonymous):

because if x is 1/5, then the denominator would be 0... and we can't divide be 0

OpenStudy (anonymous):

be careful when you simplify though, because when you bring out the x^3.. that actually does still exist. so x cannot be 0 either

OpenStudy (anonymous):

oh. ok that makes sense x) i thought there was a special way to figure it out. not just plug it in

OpenStudy (anonymous):

when we simplify, we are just reducing terms to make it easier to picture. but in this case we can factor out a x^3 from the top and bottom, which what really means is we multiply \[\frac{ 14 }{ 5x-1 }\] by \[\frac{ x^3 }{ x^3}\] which means x cannot be 0 either, because then 0 would be in the denominator of that fraction too

OpenStudy (anonymous):

so if theres an x on the bottom it cant be zero or a number that cancels out the other number for it to be 0

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

well not exactly in every case. in this example we factored out an x on the top and bottom. but lets say if we factored out \[\frac{ x^3+1 }{ x^3+1 }\] then x can still be 0

OpenStudy (anonymous):

what i meant if its connected to another number

OpenStudy (anonymous):

connect or not connect... coefficients shouldn't matter

OpenStudy (anonymous):

ok